# Simple Path Structural Encoding for Graph Transformers

### v1

@@ 1 Introduction
Graphs are pervasive across diverse domains, representing complex relationships in areas such as social networks(Otte and Rousseau, 2002), molecular structures(Quinn et al., 2017), and citation graphs(Radicchi et al., 2011). Recent advances in graph neural networks (GNNs) have driven significant progress in learning from graph-structured data (Kipf and Welling, 2017; Veličković et al., 2018; Bodnar et al., 2022; Lachi et al., 2024; Ferrini et al., 2024; Duta and Liò, 2024), yet these models often face challenges in capturing long-range dependencies and structural patterns due to their reliance on localized message passing (Alon and Yahav, 2021; Topping et al., 2022).
Inspired by their success in vision and sequence learning tasks (Vaswani, 2017; Dosovitskiy, 2020), transformers have been extended to graph learning problems (Yun et al., 2019; Dwivedi and Bresson, 2020; Ying et al., 2021; Kreuzer et al., 2021). Unlike traditional GNNs, graph transformers leverage global self-attention, allowing each node to attend to all others within a graph, regardless of distance. This flexibility overcomes the limitations of message-passing approaches but introduces new challenges, particularly in designing suitable positional and structural encodings that capture the inherent irregularities of graphs.
For directed acyclic graphs (DAGs), positional encodings (PEs) based on partial orderings can be directly applied (Dong et al., 2022; Luo et al., 2024a; Hwang et al., 2024). However, for general undirected graphs, successful PEs often rely on eigendecompositions of the graph Laplacian, drawing inspiration from sinusoidal encodings in sequence transformers (Dwivedi and Bresson, 2020; Mialon et al., 2021). These approaches encode node-level information but fail to capture the full structural complexity of edge patterns in node neighborhoods.
@@
A promising alternative is random walk structural encoding (RWSE), which encodes richer structural information by considering random walk probabilities as edge features. RWSE has shown substantial improvements in the performance of state-of-the-art graph transformers (Menegaux et al., 2023; Ma et al., 2023). However, RWSE struggles to differentiate between distinct graph structures in certain cases. Meanwhile, parallel research on message-passing GNNs has demonstrated the benefits of simple paths (or self-avoiding walks) over random walks in enhancing model expressivity (Michel et al., 2023; Graziani et al., 2023). These findings motivate the exploration of simple paths as a structural encoding mechanism in graph transformers.
In this work, we introduce Simple Path Structural Encoding (SPSE), a novel method for structural edge encoding that replaces RWSE in graph transformers. SPSE encodes graph structure by counting simple paths of varying lengths between node pairs, capturing richer structural information than random walks. To address the computational challenges of simple path counting, we propose an efficient algorithm based on successive DAG decompositions using depth-first search (DFS) and breadth-first search (BFS). This approach avoids the exponential memory costs of path enumeration, enabling scalability to long path lengths.  [^1] The Python code for the path counting algorithm is available on the project’s page: https://github.com/LouisBearing/Graph-SPSE-Encoding.
@@ 2.1 Graph Theory
Similarly, we define the $k$-hop simple path matrix $S_{k}\in\mathbb{N}^{|\mathcal{V}|\times|\mathcal{V}|}$ as the matrix where the ($i,j$)-th entry, $(S_{k})_{ij}$, is the number of $k$-length simple paths from node $i$ to node $j$.
Unlike $P_{k}$, $S_{k}$ does not have a closed-form solution, and computing it is computationally expensive (Vassilevska and Williams, 2009). In the following, we focus on the properties of pairs of nodes $(i,j)\in\mathcal{V}\times\mathcal{V}$ through different edge encoding methodologies. To facilitate the comparison with the encoding of another pair of nodes $(i^{\prime},j^{\prime})$ in a second graph $\mathcal{G^{\prime}}$, we introduce the following equivalence relation for $k\geq 1$:
@@ 3 Theoretical Properties of RWSE and SPSE
In this section, we highlight the limitations of RWSE in distinguishing different graph structures, as it can assign identical transition probabilities to edges in distinct topologies, potentially leading to suboptimal performance. Conversely, we show that SPSE naturally captures cycle-related information. In particular, SPSE enables cycle counting, which is crucial in applications such as molecular chemistry (e.g., identifying functional groups like aromatic rings) (May and Steinbeck, 2014; Agúndez et al., 2023), social network analysis (e.g., detecting communities) (Radicchi et al., 2004; Dhilber and Bhavani, 2020), and circuit design (e.g., analyzing feedback loops) (Horowitz et al., 1989)
@@ 3.1 Limitations of RWSE for Edge Encoding
Figure 1: RWSE encodes identically edges from very distinct graphs which are separated by SPSE (see the tree decompositions in Appendix C)
Random walk probabilities possess the great advantage of being computable in closed form. However, their use introduces certain ambiguities, as illustrated in Figure 1. In these two examples, $\mathcal{G}_{A}$ and $\mathcal{G}_{C}$ are cycle graphs, $\mathcal{G}_{B}$ and $\mathcal{G}_{D}$ are path graphs, and yet the following two relations hold (see Appendix C for a graphical proof up to a depth of 5):
@@
Although the edges belong to clearly different graphs, RWSE assigns them the same transition probabilities in the cycle and path graphs. These examples are in reality two special cases of the following result which links RWSE edge encodings in even-length cycle graphs and linear graphs (all proofs can be found in Appendix B):
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an even-length cycle graph, i.e. $|\mathcal{V}|=2n$ for some $n\in\mathbb{N}^{*}$, and $\mathcal{G^{\prime}}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime})$ a path graph such that $|\mathcal{V}^{\prime}|=2n+1$. Then given any pair of nodes $(i,j)$ in $\mathcal{G}$, there exits a pair of nodes $(i^{\prime},j^{\prime})$ in $\mathcal{G^{\prime}}$ such that $(i,j)^{\mathcal{G}}=_{\textsc{rw}}(i^{\prime},j^{\prime})^{\mathcal{G^{\prime}}}$.
Each pair of nodes in an even-length cycle graph is thus equivalent, under RWSE encoding, to another pair of nodes in a linear graph. In other words, random walks transition probabilities cannot be used to distinguish between even-length cycles and paths when considering single node pairs. On the other end, it is easy to show that this does not apply to SPSE encoding (see Figure 1). This result illustrates how random walk-based structural encoding may fail to capture critical structural differences between graphs, hence possibly leading to suboptimal performances of the overall graph transformer model.
@@
The RWSE encoding of a pair of nodes (note that $i$ and $j$ need not be adjacent) is therefore never unique, and cannot be used to identify a graph. Of course this is also obviously true for SPSE encoding. However, proving this result for adjacent nodes, in the cases of simple paths, still requires preserving the set of paths that connect them. This hints at the fact that the information contained in simple path edge encodings may be used to distinguish certain local graph structures, as we discuss in the next section.
@@ 3.2 SPSE through the Prism of Cycle Counting
The relation between path and cycle counting has been well studied (Perepechko and Voropaev, 2009; Graziani et al., 2023). The following result, which was introduced by Perepechko and Voropaev (2009), connects SPSE encoding with cycle counting for adjacent nodes:
@@ 4 Simple Path Structural Encoding
The superior expressivity of paths over walks in MPNNs (Michel et al., 2023; Graziani et al., 2023) has been studied, where message-passing follows short paths stored in memory. However, due to the combinatorial complexity of enumerating all paths, storing these sequences becomes impractical. Results from Section 3 indicate that counting distinct paths between nodes, while requiring significantly less memory, still possesses theoretical advantages over random walk probabilities as an edge structural encoding method. This approach, however, introduces two challenges. First, while existing path-counting algorithms efficiently handle short paths (Perepechko and Voropaev, 2009; Giscard et al., 2019), the graph topologies and path lengths considered here (15-23) necessitate approximate methods. Second, since the number of paths between two nodes can grow exponentially (bounded by $\frac{(|\mathcal{V}|-2)!}{(|\mathcal{V}|-k-1)!}$ for length-$k$ paths in a complete graph), an appropriate encoding function is required. This section addresses both challenges.
@@ 4.1 Simple Path Counting
The path counting method followed here relies on the extraction of node orderings from an input undirected graph. Any such ordering can be used to turn the graph into a DAG, which allows to count paths by computing powers of the resulting adjacency matrix. For paths of length $K$, this process incurs a computational complexity of $\mathcal{O}(K|\mathcal{V}|^{3})$. Different DAG decompositions of the input undirected graph allow the exploration of different paths, leading to increasingly more accurate path counts. We summarize the overall procedure in Algorithm 1. The detailed version of this algorithm can be found in Appendix D. It consists in updating running path counts for all pairs of nodes and path length $k$, by comparing the counts yielded by each DAG with the stored values. The DAG mining function is presented in Algorithm 2. Starting from a root node, it decomposes a graph into a tree by combining DFS and BFS: a DFS search is initiated at a given root node, updating a list which is incremented with each newly visited node. After it reaches a distance $D_{\textsc{dfs}}$ from the root node, the search switches to BFS until it cannot proceed further, and the whole process is repeated until all nodes are visited. A DAG can be obtained from the node ordering by directing edges towards nodes of higher indices. DFS allows long path discoveries but cannot simultaneously discover more than one path between nodes, while BFS alone will typically miss long paths, hence their use in combination. We add a partial BFS step between the two, which randomly explores a subset of the child nodes, thus allowing to travel through otherwise inaccessible paths (see Figure 3 for an illustration). Additional parameters include a repetition number $N$ which accounts for random effects, such as the direction taken by the DFS, and a parameter $R$ controlling the proportion of root nodes. The computational cost is dominated by the powers of the adjacency matrix which must be computed for each DAG, for a maximal complexity of $\mathcal{O}(KRD_{\textsc{dfs}}N|\mathcal{V}|^{3})$. The additional factor $RD_{\textsc{dfs}}N$ can be of the order of tens to hundreds depending on the input graph, which makes SPSE calculation significantly more expensive than RWSE. This needs however to be computed only once as a pre-processing step, and it is also much less than the total number of possible DAG decompositions $2^{|\mathcal{E}|}$, highlighting the effectiveness of the tree decomposition approach. We discuss the choice of hyperparameters and their importance regarding path counts in Section 5.3 and provide actual values in Appendix A.
@@ 5.1 Cycle Counting Synthetic Experiment
To validate Proposition 3, we design a synthetic dataset consisting of 12,000 graphs. Each graph is generated by randomly adding cycles of lengths between 3 and 8 until the total count for each cycle length reaches a value between 0 and 14. This process results in graphs with an average of 149 nodes and 190 edges. The dataset is split into training (10,000 graphs), validation (1,000 graphs), and test (1,000 graphs) sets. The objective is to determine the number of cycles for each of the six cycle lengths. We frame this as six simultaneous multiclass classification tasks and evaluate performance using mean accuracy. Examples of the generated graphs are provided in Appendix E.
@@
Table 1: Test results on all eight benchmarks. GPS, CSA & GRIT w/ and w/o SPSE are (re-)trained on 10 random seeds (hence variations from seminal works) on all datasets (one run for PCQM4Mv2). Underline indicates a difference between a baseline and the SPSE model version, and * is used for significant gaps based on a two-sided t-test ($p$-value $\leq 0.05$). Highlighted are best, second, and third best scores.
@@
CSA-RWSE $0.069{\pm 0.003}$ $0.6513\pm 0.0061$ $0.24862\pm 0.0012$ $0.0918$ $87.008{\pm 0.062}$ $79.071{\pm 0.120}$ $98.127{\pm 0.123}$ $73.885{\pm 0.348}$
CSA-SPSE (ours) $\textbf{0.061}{\pm\textbf{0.003}}$* $0.6605\pm 0.0096$* $0.24945\pm 0.0018$ $0.0911$ $\textbf{87.064}{\pm\textbf{0.052}}$* $78.940{\pm 0.132}$ ${\color[rgb]{0.55,0,0}\underline{\textbf{98.269}{\pm\textbf{0.078}}}}$* $73.897{\pm 0.524}$
GRIT-RWSE ${\color[rgb]{0.55,0,0}\textbf{0.065}{\pm\textbf{0.005}}}$ $0.6803{\pm 0.0085}$ $\textbf{0.2480}{\pm\textbf{0.0025}}$ 0.0838 $\textbf{87.229}{\pm\textbf{0.056}}$ ${\color[rgb]{0,1,1}\underline{\textbf{79.730}{\pm\textbf{0.189}}}}$ $98.231{\pm 0.197}$ $\textbf{76.246}{\pm\textbf{0.954}}$
GRIT-SPSE (ours) ${\color[rgb]{0,1,1}\underline{\textbf{0.059}{\pm\textbf{0.001}}}}$* $\textbf{0.6945}{\pm\textbf{0.0113}}$* ${\color[rgb]{0,1,1}\underline{\textbf{0.2449}{\pm\textbf{0.0018}}}}$* ${\color[rgb]{0,1,1}\underline{\textbf{0.0831}}}$ ${\color[rgb]{0,1,1}\underline{\textbf{87.235}{\pm\textbf{0.040}}}}$ $\textbf{79.571}{\pm\textbf{0.122}}$ $\textbf{98.294}{\pm\textbf{0.147}}$ ${\color[rgb]{0,1,1}\underline{\textbf{77.022}{\pm\textbf{0.430}}}}$*
Results.
The test accuracy for both model architectures across the three training configurations is reported in Figure 4, along with standard deviations computed over 10 runs with different random seeds. In all but one case, SPSE encoding achieves significantly higher cycle counting accuracy than RWSE. All models learn to count cycles almost perfectly under the third training configuration, which suggests that deep architectures can compensate for expressivity limitations in the edge encoding matrix (see Appendix E for details on configurations). However, SPSE still performs significantly better than RWSE for CSA in this setting.

### v2

@@ Abstract
Graph neural networks, transformer, structural encoding, simple paths
@@ 1 Introduction
Graphs are pervasive across diverse domains, representing complex relationships in areas such as social networks (Otte & Rousseau, 2002), molecular structures (Quinn et al., 2017), and citation graphs (Radicchi et al., 2011). Recent advances in graph neural networks (GNNs) have driven significant progress in learning from graph-structured data (Kipf & Welling, 2017; Veličković et al., 2018; Bodnar et al., 2022; Lachi et al., 2024; Ferrini et al., 2024; Duta & Liò, 2024), yet these models often face challenges in capturing long-range dependencies and structural patterns due to their reliance on localized message passing (Alon & Yahav, 2021; Topping et al., 2022).
Inspired by their success in vision and sequence learning tasks (Vaswani, 2017; Dosovitskiy et al., 2021), transformers have been extended to graph learning problems (Yun et al., 2019; Dwivedi & Bresson, 2020; Ying et al., 2021; Kreuzer et al., 2021). Unlike traditional GNNs, graph transformers leverage global self-attention, allowing each node to attend to all others within a graph, regardless of distance. This flexibility overcomes the limitations of message-passing approaches but introduces new challenges, particularly in designing suitable positional and structural encodings that capture the inherent irregularities of graphs.
For directed acyclic graphs (DAGs), positional encodings (PEs) based on partial orderings can be directly applied (Dong et al., 2022; Luo et al., 2024b; Hwang et al., 2024). However, for general undirected graphs, successful PEs often rely on eigendecompositions of the graph Laplacian, drawing inspiration from sinusoidal encodings in sequence transformers (Dwivedi & Bresson, 2020; Mialon et al., 2021). These approaches encode node-level information but fail to capture the full structural complexity of edge patterns in node neighborhoods.
@@
A promising alternative is random walk structural encoding (RWSE), which encodes richer structural information by considering random walk probabilities as edge features. RWSE has shown substantial improvements in the performance of state-of-the-art graph transformers (Menegaux et al., 2023; Ma et al., 2023). However, RWSE struggles to differentiate between distinct graph structures in certain cases. Meanwhile, parallel research on message-passing GNNs has demonstrated the benefits of simple paths (or self-avoiding walks) in enhancing model expressivity beyond the 1-Weisfeiler-Leman (WL) isomorphism test (Michel et al., 2023; Graziani et al., 2024). These findings motivate the exploration of simple paths as a structural encoding mechanism in graph transformers.
In this work, we introduce Simple Path Structural Encoding (SPSE), a novel method for structural edge encoding that replaces RWSE in graph transformers. SPSE encodes graph structure by counting simple paths of varying lengths between node pairs, capturing richer structural information than random walks. To address the computational challenges encountered when exact simple path counting is unfeasible, we propose an efficient algorithm based on successive DAG decompositions using depth-first search (DFS) and breadth-first search (BFS). This approach avoids the exponential memory costs of path enumeration, enabling scalability to long path lengths.  [^1] The Python implementation of the algorithm is available on the project’s Github page.
@@ 2.1 Graph Theory
Similarly, we refer to the $k$-hop simple path matrix $S_{k}\in\mathbb{N}^{|\mathcal{V}|\times|\mathcal{V}|}$ as the matrix where the ($i,j$)-th entry, $(S_{k})_{ij}$, is the number of simple paths of length $k$ from node $i$ to node $j$.
Unlike $P_{k}$, $S_{k}$ does not have a closed-form solution, and computing it is computationally expensive (Vassilevska & Williams, 2009). In the following, we focus on the properties of pairs of nodes $(i,j)\in\mathcal{V}\times\mathcal{V}$ through different edge encoding methodologies. To facilitate the comparison with the encoding of another pair of nodes $(i^{\prime},j^{\prime})$ in a second graph $\mathcal{G^{\prime}}$, we introduce the following equivalence relation for $k\geq 1$:
@@ 3 Theoretical Properties of RWSE and SPSE
In this section, we highlight the limitations of RWSE in distinguishing different graph structures, as it can assign identical transition probabilities to edges in distinct topologies, potentially leading to suboptimal performance. Conversely, we show that SPSE naturally captures cycle-related information. In particular, SPSE enables cycle counting, which is crucial in applications such as molecular chemistry (e.g., identifying functional groups like aromatic rings) (May & Steinbeck, 2014; Agúndez et al., 2023), social network analysis (e.g., detecting communities) (Radicchi et al., 2004; Dhilber & Bhavani, 2020), and circuit design (e.g., analyzing feedback loops) (Horowitz et al., 1989).
@@ 3.1 Limitations of RWSE for Edge Encoding
Figure 1: RWSE encodes identically edges from very distinct graphs which are separated by SPSE (see the tree decompositions in Appendix D)
Random walk probabilities possess the great advantage of being computable in closed form. However, their use introduces certain ambiguities, as illustrated in Figure 1. In these two examples, $\mathcal{G}_{A}$ and $\mathcal{G}_{C}$ are cycle graphs, $\mathcal{G}_{B}$ and $\mathcal{G}_{D}$ are path graphs, and yet the following two relations hold (see Appendix D for a graphical proof up to a depth of 5):
@@
Although the edges belong to clearly different graphs, RWSE assigns them the same transition probabilities in the cycle and path graphs. These examples are two special cases of the following result which links RWSE edge encodings in even-length cycle graphs and linear graphs (all proofs can be found in Appendix C):
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an even-length cycle graph, i.e. $|\mathcal{V}|=2n$ for some $n\in\mathbb{N}^{*}$, and $\mathcal{G^{\prime}}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime})$ a path graph such that $|\mathcal{V}^{\prime}|=2n+1$. Then given any pair of nodes $(i,j)$ in $\mathcal{G}$, there exists a pair of nodes $(i^{\prime},j^{\prime})$ in $\mathcal{G^{\prime}}$ such that $(i,j)^{\mathcal{G}}=_{\textsc{rw}}(i^{\prime},j^{\prime})^{\mathcal{G^{\prime}}}$.
Each pair of nodes in an even-length cycle graph is thus equivalent, under RWSE encoding, to another pair of nodes in a linear graph (note that $i$ and $j$ need not be adjacent). In other words, random walks transition probabilities cannot be used to distinguish between even-length cycles and paths when considering single node pairs. On the other end, it is easy to show that this does not apply to SPSE encoding (see Figure 1). This result illustrates how random walk-based structural encoding may fail to capture critical structural differences between node pairs, hence possibly leading to suboptimal performances of the overall graph transformer model.
@@
The RWSE encoding of a pair of nodes is therefore never unique, and cannot be used to identify a graph. Of course this is also obviously true for SPSE encoding. However, proving this result for adjacent nodes, in the cases of simple paths, still requires preserving the set of paths that connect them. This hints at the fact that the information contained in simple path edge encodings may be used to distinguish certain local graph structures, as we discuss in the next section.
@@ 3.2 SPSE through the Prism of Cycle Counting
The relation between path and cycle counting has been well studied (Perepechko & Voropaev, 2009; Graziani et al., 2024). The following result, which was introduced by Perepechko & Voropaev (2009), connects SPSE encoding with cycle counting for adjacent nodes:
@@
Finally, while the results presented here focus on edge representations, we refer to Appendix B for a discussion on the expressivity of SPSE regarding graph isomorphism.
@@ 4 Simple Path Structural Encoding
It has been shown that the composition of MPNN layers on paths is more expressive than the 1-WL test (Michel et al., 2023; Graziani et al., 2024). However, due to the combinatorial complexity of the problem, enumerating all paths beyond short path lengths becomes impractical. Results from Section 3 indicate that counting distinct paths between nodes, while requiring significantly less memory, still possesses theoretical advantages over random walk probabilities as an edge structural encoding method. This approach, however, introduces two challenges. First, while existing path-counting algorithms efficiently handle short paths (Perepechko & Voropaev, 2009; Giscard et al., 2019), certain graph topologies and path lengths necessitate approximate methods. Second, since the number of paths between two nodes can grow exponentially (bounded by $\frac{(|\mathcal{V}|-2)!}{(|\mathcal{V}|-k-1)!}$ for length-$k$ paths in a complete graph), an appropriate encoding function is required. This section addresses both challenges.
@@ 4.1 Simple Path Counting
The SPSE encoding can accommodate any exact or approximate path counting method, although the former are usually restricted to sparse graphs or short path lengths. On the other hand, approximate path counting methods are required when the density of the considered graphs prevents the use of exact methods.
The approximate method followed here relies on the extraction of node orderings from an input undirected graph. Any such ordering can be used to turn the graph into a DAG, which allows to count paths by computing powers of the resulting adjacency matrix. For paths of length $K$, this process incurs a computational complexity of $\mathcal{O}(K|\mathcal{V}|^{3})$. Different DAG decompositions of the input undirected graph allow the exploration of different paths, leading to increasingly more accurate path counts. We summarize the overall procedure in Algorithm 1. The detailed version of this algorithm can be found in Appendix E. It consists in updating running path counts for all pairs of nodes and path length $k$, by comparing the counts yielded by each DAG with the stored values, and then storing the maximum of the two. The DAG mining function is presented in Algorithm 2. Starting from a root node, it decomposes a graph into a tree by combining DFS and BFS: a DFS search is initiated at a given root node, updating a list which is incremented with each newly visited node. After it reaches a distance $D_{\textsc{dfs}}$ from the root node, the search switches to BFS until it cannot proceed further, and the whole process is repeated until all nodes are visited. A DAG can be obtained from the node ordering by directing edges towards nodes of higher indices. DFS allows long path discoveries but cannot simultaneously discover more than one path between nodes, while BFS alone will typically miss long paths, hence their use in combination. We add a partial BFS step between the two, which randomly explores a subset of the child nodes, thus allowing to travel through otherwise inaccessible paths (see Figure 3 for an illustration). Additional parameters include a repetition number $N$ which accounts for random effects, such as the direction taken by the DFS, and a parameter $R$ controlling the proportion of root nodes. The computational cost is dominated by the powers of the adjacency matrix which must be computed for each DAG, for a maximal complexity of $\mathcal{O}(KRD_{\textsc{dfs}}N|\mathcal{V}|^{3})$. The additional factor $RD_{\textsc{dfs}}N$ can be of the order of tens to hundreds depending on the input graph, which makes SPSE calculation significantly more expensive than RWSE. This needs however to be computed only once as a pre-processing step, and it is also much less than the total number of possible DAG decompositions $2^{|\mathcal{E}|}$, highlighting the effectiveness of the tree decomposition approach. We discuss the choice of hyperparameters and their importance regarding path counts in Section 5.3 and provide actual values in Appendix A.
@@ 5.1 Cycle Counting Synthetic Experiment
To validate Proposition 3, we design a synthetic dataset consisting of 12,000 graphs. Each graph is generated by randomly adding cycles of lengths between 3 and 8 until the total count for each cycle length reaches a value between 0 and 14. This process results in graphs with an average of 149 nodes and 190 edges. The dataset is split into training (10,000 graphs), validation (1,000 graphs), and test (1,000 graphs) sets. The objective is to determine the number of cycles for each of the six cycle lengths. We frame this as six simultaneous multiclass classification tasks and evaluate performance using mean accuracy. Examples of the generated graphs are provided in Appendix F.
@@
Results.
The test accuracy for both model architectures across the three training configurations is reported in Figure 4, along with standard deviations computed over 10 runs with different random seeds.
In all but one case, SPSE encoding achieves significantly higher cycle counting accuracy than RWSE. All models learn to count cycles almost perfectly under the third training configuration, which suggests that deep architectures can compensate for expressivity limitations in the edge encoding matrix (see Appendix F for details on configurations). However, SPSE still performs significantly better than RWSE for CSA in this setting.
@@ Results.
Table 1: Test results on all eight benchmarks. GPS, CSA & GRIT w/ and w/o SPSE were re-trained on 10 random seeds (hence variations from seminal works) on all datasets (one run for PCQM4Mv2). Underline indicates a difference between a baseline and the SPSE model version, and * is used for significant gaps based on a two-sided t-test ($p$-value $\leq 0.05$). Highlighted are best, second, and third best scores. GPS+GPSE was not retrained and is therefore treated separately.
@@
CSA-RWSE $0.069{\pm 0.003}$ $0.6513\pm 0.0061$ $0.2486\pm 0.0012$ $0.0918$ $87.008{\pm 0.062}$ $79.071{\pm 0.120}$ $98.127{\pm 0.123}$ $73.885{\pm 0.348}$
CSA-SPSE (ours) $\textbf{0.061}{\pm\textbf{0.003}}$* $0.6605\pm 0.0096$* $0.2482\pm 0.0019$ $0.0911$ $\textbf{87.064}{\pm\textbf{0.052}}$* $78.940{\pm 0.132}$ ${\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\underline{\textbf{98.269}{\pm\textbf{0.078}}}}$* $73.897{\pm 0.524}$
GRIT-RWSE ${\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\textbf{0.065}{\pm\textbf{0.005}}}$ $0.6803{\pm 0.0085}$ $0.2480\pm 0.0025$ 0.0838 $\textbf{87.229}{\pm\textbf{0.056}}$ ${\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\underline{\textbf{79.730}{\pm\textbf{0.189}}}}$ $98.231{\pm 0.197}$ $\textbf{76.246}{\pm\textbf{0.954}}$
GRIT-SPSE (ours) ${\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\underline{\textbf{0.059}{\pm\textbf{0.001}}}}$* $\textbf{0.6945}{\pm\textbf{0.0113}}$* ${\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\underline{\textbf{0.2449}{\pm\textbf{0.0018}}}}$* ${\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\underline{\textbf{0.0831}}}$ ${\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\underline{\textbf{87.235}{\pm\textbf{0.040}}}}$ $\textbf{79.571}{\pm\textbf{0.122}}$ $\textbf{98.294}{\pm\textbf{0.147}}$ ${\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\underline{\textbf{77.022}{\pm\textbf{0.430}}}}$*